OpenStudy (anonymous):

A cubical block of metal weighs 6 pounds. How much will another cube of the same metal weigh if its sides are twice as long?

OpenStudy (anonymous):

you have any ideas?

pooja195 (pooja195):

assume that the cubical block's sides measures are all x: volume = x*x*x = x³ if the sides measures become doubled, each side will become 2x: volume = 2x*2x*2x = 8x³ so the volume will increase 8 times if the sides are doubled, and since volume is directly proportional to weight ---> weight increases by 8 times: 6*8 = 48 pounds